OpenStudy (anonymous):

the directional derivative of f(xyz) = xy^3z^2 at point p(2,-1.4), in the direction of vector a= <1,2,-1> is 192/root(6) the vector is -16/root(6), 192/root(6), 16/root(6) in what direction does f increase most rapidly from the point p? (the answer is a unit vector)

OpenStudy (anonymous):

\[Du|_f = 192/\sqrt(6)\] \[<-16/\sqrt(6), 192 / \sqrt(6), 16/ \sqrt(6)>\]

OpenStudy (anonymous):

the most rapidly increas direction for a function is it's gradient direction. So find the grad f.

OpenStudy (anonymous):

i'm pretty sure the gradient is the vector above, no?

OpenStudy (anonymous):

\[grad f=(y ^{3}z ^{2},3xy ^{2}z ^{2},2xy ^{3}z)\] at point p(2,-1,4) grad f =(-16,96,-16) check just in case, maybe i made some calculation error....

OpenStudy (anonymous):

is the gradient: \[Deltaf= (y^3x^2)i+ (3y^2xz^3)j+(2zxy^3)k\]

OpenStudy (anonymous):

should be z^2 in the j component

OpenStudy (anonymous):

and z^2 instead of x^2 in the i one

OpenStudy (anonymous):

it will increase most rapidly at the direction which is parallel to our gradient vector. now our gradient vector is just <df/dx,df/dy,df/dz>=<y^3z^2,3xy^2z^2,2xy^3z>, now at point (2,-1,4), well have: =<-16,96,-16> now we just need to turn this to a unit vector which is parallel to the gradient vector, so we just multiply by the reciprocal of the norm of this gradient. so the norm is just: =sqrt(9728)=16sqrt(38) so the answer will be: =<-16/16sqrt(38),96/16sqrt(38),16/sqrt(38)>=<-1/sqrt(38),6/sqrt(38),1/sqrt(38)> this should be the direction of our unit vector

OpenStudy (anonymous):

you lost me around here: now we just need to turn this to a unit vector which is parallel to the gradient vector, so we just multiply by the reciprocal of the norm of this gradient. so the norm is just: =sqrt(9728)=16sqrt(38) so the answer will be: =<-16/16sqrt(38),96/16sqrt(38),16/sqrt(38)>=<-1/sqrt(38),6/sqrt(38),1/sqrt(38)> this should be the direction of our unit vector

OpenStudy (anonymous):

at point p(2,-1,4) grad f =(-16,96,-16) this is not unit vector. To make it unit vector you have to devide it by it's length. The length of the vector v=(x,y,z) is:\[|v|=\sqrt{x ^{2}+y ^{2}+z ^{2}}\]

OpenStudy (anonymous):

in your case the length is:\[length=\sqrt{(-16)^{2}+96^{2}+(-16)^{2}}\]

OpenStudy (anonymous):

how is that different from the first vector i gave you?

OpenStudy (anonymous):

i'm really confused

OpenStudy (anonymous):

if its in the direction of a = <1,2,-1> then isn't the unit vector \[1/\sqrt(6) <-16, 96, -16>\]

OpenStudy (anonymous):

it is not in the direction of a

OpenStudy (anonymous):

it is in the direction of grad f

OpenStudy (anonymous):

wait, no that's wrong i took the gradient <-16, 96, -16> and dotted it with the unit vector of a to get \[1/\sqrt(6) <-16,192,-16>\]

OpenStudy (anonymous):

so i no longer care about a?

OpenStudy (anonymous):

this way you get the derivative in the direction of vector a, but you were asking about the direction where f increas is maximum. And this maximum increas direction is direction of the grad f

OpenStudy (anonymous):

ohhhhhhhh so i have to get the unit vector of f ohhh okay i think it's starting to make sense haha

OpenStudy (anonymous):

unit of grad f

OpenStudy (anonymous):

so the unit vector of = \[1/\sqrt(9728) <-16, 96, -16>\] ?

OpenStudy (anonymous):

yes, you just can simplify this a bit

OpenStudy (anonymous):

i'm not sure how to simplify it /blush

OpenStudy (anonymous):

oh.. 16x16.. ohh.... hahha

OpenStudy (anonymous):

9728=256*38=16*16*38

OpenStudy (anonymous):

so it's \[1/\sqrt(38) <-1, 6, -1>\]?

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

yayyyy thank you :D :D :D